OpenStudy (anonymous):

2x+2y=19 and 2x-2y=5 solve using the addition method

OpenStudy (anonymous):

plzzzz someone help me i need to get my work done tonight plzzzzzzzzzz

OpenStudy (gw2011):

2x+2y=19 2x-2y=5 Add these equations and you get: 4x=24 x=6 Substitute x=6 into 2x+2y=19 (2)(6)+2y=19 12+2y=19 2y=7 y=7/2 Therefore, x=6 and y=7/2

OpenStudy (anonymous):

2x-2y=5 .. Add 14 to both sides ... 2x-2y+14 =19... We know 2x+2y=19 .. so substituting that we get 2x-2y+14 = 2x+2y .. 2x get cancelled out we get y =14/4 and x as 24/4 ... this can simply be solved adding two equations also.

OpenStudy (anonymous):

thanks i got another 1 using the addition methd 2x-y=8 and 4x+3y=36

OpenStudy (gw2011):

4x+3y=36 2x-y=8 Multiply this equation by 2 and you get: 4x+3y=36 4x-2y=16 Subtract this equation from the above equation and you get: 5y=20 y=4 Substitute y=4 into 2x-y=8: 2x-y=8 2x-4=8 2x=12 x=6 Therefore, x=6 and y=4